# Balancing Quality and Variation: Spam Filtering Distorts Data Label Distributions

### v1

@@ Appendix A Computing the CrowdTruth Worker Quality Score
As highlighted in Section 4.2, WWA measures how similar a given worker’s annotations are to other workers, weighted by the workers’ quality and the instances’ (or units’) ambiguity. WUA measures how much a worker agrees with the aggregate label over all their annotated instances, weighted by the instances’ ambiguity. WWA and WUA are roughly computed as follows (ignoring the normalization terms for clarity, full details in Dumitrache et al., 2018b):
@@ Appendix D Dataset Selection Table
DICES Aroyo et al. 2023 Yes Yes
MTurk Survey Huang et al. 2023 Yes Yes
MHS corpus Sachdeva et al. 2022 Yes No Raters excluded (details in their paper), but data not available.
AdultContent3 (”Get Another Label“ datasets) Ipeirotis et al. 2010 No No No gold spammers. Experiments in paper use synthetic data and simply report impact on the collected dataset
HITspam Discussed in Ertekin et al. 2014 No No Despite the name, does not contain spammers. Instead, the task is to judge whether a task on MTurk itself should be considered spam (e.g., because it asks workers to follow a specific social media account).
EDOS-DOM Jiang et al. 2024 Yes No Only one annotator removed after labels were collected. That annotator had annotated only 8 examples (first author vial email).
Argument Quality Mirzakhmedova et al. 2024 No No Excludes a number of disagreeing annotations per example. Does not exclude on the level of the annotator.
MultiPref Miranda et al. 2025 No No No gold spammers.
HelpSteer2 Wang et al. 2025 No No No gold spammers.
CrowdTruth Corpus for Open Domain Relation Extraction Dumitrache et al. 2017 No No Emailed first author, full data not available anymore.
AMR / Sentence Similarity Data Wein and Schneider 2022 Yes No Only one annotator removed out of three in total.
Phrase Detectives Chamberlain et al. 2016 Yes, self-reported No Spammer data not available anymore according to authors.
Crowd-Sourced Preference Tests Buchholz and Latorre 2011 Yes, inferred No Data not available anymore according to first author.
VariErr NLI Weber-Genzel et al. 2024 No No Data has annotator IDs and individual decisions plus error judgments (error = no self-validations), but no excluded raters. Not a crowd-sourced study (four annotators).
Malicious Worker Survey Dataset Gadiraju et al. 2015 Yes No Dataset is not available anymore.
Dog (Imagenet Subset) Deng et al. 2009 No No No spammer information. Used by Traganitis and Giannakis (2021), but only evaluated by accuracy of resulting classifier.
ImageNetV2 Recht et al. 2019 No No No spammer information.
Bluebird Welinder et al. 2010 No No No spammer information. Used by Traganitis and Giannakis (2021), but only evaluated by accuracy of resulting classifier.
Web Zhou et al. 2012 Unlikely No Could not find reference to data.
WSD Snow et al. 2008 Unlikely No Data not available anymore
RTE Snow et al. 2008 Unlikely No Data not available anymore
TEMP Snow et al. 2008 Unlikely No Data not available anymore
POPQUORN Pei and Jurgens 2023 Yes No Only a single annotator was removed.

### v2

@@ Appendix A Computing the CrowdTruth Worker Quality Score
As highlighted in Section 4.2, WWA measures how similar a given worker’s annotations are to other workers, weighted by the workers’ quality and the instances’ (or units’) ambiguity. WUA measures how much a worker agrees with the aggregate label over all their annotated instances, weighted by the instances’ ambiguity. WWA and WUA are roughly computed as follows (ignoring the normalization terms for clarity, full details in CrowdTruth2):
@@ Appendix D Dataset Selection Table
DICES aroyo2023dices Yes Yes
MTurk Survey huang-etal-2023-incorporating Yes Yes
MHS corpus sachdeva_measuring_2022 Yes No Raters excluded (details in their paper), but data not available.
AdultContent3 (”Get Another Label“ datasets) ipeirotis_quality_2010 No No No gold spammers. Experiments in paper use synthetic data and simply report impact on the collected dataset
HITspam Discussed in ertekin_approximating_2014 No No Despite the name, does not contain spammers. Instead, the task is to judge whether a task on MTurk itself should be considered spam (e.g., because it asks workers to follow a specific social media account).
EDOS-DOM jiang-etal-2024-examining Yes No Only one annotator removed after labels were collected. That annotator had annotated only 8 examples (first author vial email).
Argument Quality mirzakhmedova_are_2024 No No Excludes a number of disagreeing annotations per example. Does not exclude on the level of the annotator.
MultiPref miranda-etal-2025-hybrid No No No gold spammers.
HelpSteer2 wang_helpsteer2_2025 No No No gold spammers.
CrowdTruth Corpus for Open Domain Relation Extraction dumitrache_false_2017 No No Emailed first author, full data not available anymore.
AMR / Sentence Similarity Data wein-schneider-2022-accounting Yes No Only one annotator removed out of three in total.
Phrase Detectives chamberlain_phrase_2016 Yes, self-reported No Spammer data not available anymore according to authors.
Crowd-Sourced Preference Tests buchholz_crowdsourcing_2011 Yes, inferred No Data not available anymore according to first author.
VariErr NLI weber-genzel-etal-2024-varierr No No Data has annotator IDs and individual decisions plus error judgments (error = no self-validations), but no excluded raters. Not a crowd-sourced study (four annotators).
Malicious Worker Survey Dataset gadiraju_understanding_2015 Yes No Dataset is not available anymore.
Dog (Imagenet Subset) 5206848 No No No spammer information. Used by traganitis_identifying_2021, but only evaluated by accuracy of resulting classifier.
ImageNetV2 pmlr-v97-recht19a No No No spammer information.
Bluebird welinder_multidimensional_2010 No No No spammer information. Used by traganitis_identifying_2021, but only evaluated by accuracy of resulting classifier.
Web zhou_learning_2012 Unlikely No Could not find reference to data.
WSD snow-etal-2008-cheap Unlikely No Data not available anymore
RTE snow-etal-2008-cheap Unlikely No Data not available anymore
TEMP snow-etal-2008-cheap Unlikely No Data not available anymore
POPQUORN pei-jurgens-2023-annotator Yes No Only a single annotator was removed.